OpenStudy (anonymous):

Make a 5% offer on a Volkswagen New Beetle that has a sticker price of $22,120 and a dealer cost of $20,853.

OpenStudy (anonymous):

I am willing to pay 1 dollar per question if you send me your paypal account, i need like 5 questions asap. anyone?

OpenStudy (anonymous):

I don't know what a 5% offer is, does your book or class provide any details?

OpenStudy (anonymous):

Thats why i ams o confuse, this is why i got so far, but i dont know what will be the answer if is the $1,106 or the $148. This is my calculations 5 % of 22,120=22,120*5/100=1,106 so price after of offer= 22,120-1,106 = $21,014 now dealer cost=$ 20,866 offer profit=21,014-20,866 =$148